Q 3.29 Pheromones are a special type of compound secreted by the females of many insect species to attract the males for mating. One pheromone has the molecular formula C19H38O. Normally, the amount of this pheromone secreted by a female insect is about 1.0× 10-12 g. How many molecules are there in this quantity?


1
Expert's answer
2021-10-22T03:57:41-0400

Let us first calculate the molar mass of C19H38O knowing the fact that the weight of one carbon atom is 12.01 g, one hydrogen atom is 1.008 g, and one oxygen atom is 16 g.

MM(C19H38O) "= 19 \times 12.01 + 38 \times 1.008 + 1 \times 16 = 282.5 \;g/mol"

The molar mass is the weight of 1 mole of the compound. Using this information we can find the number of moles in "1.0 \times 10^{-12} \;g" of C19H38O.

"n =\frac{1.0 \times 10^{-12}}{282.5} = 3.5 \times 10^{-15} \;mol"

One mole of a substance contains Avogadro number of molecules. Then, the number of molecules in "3.5 \times 10^{-15} \;mol" of the substance will be

"6.022 \times 10^{23} \times 3.5 \times 10^{-15} = 2.1 \times 10^9 \; molecules"

Thus, there are "2.1 \times 10^9 \; molecules" in "1.0 \times 10^{-12} \;g" of C19H38O.